OpenStudy (anonymous):

Find the quotient and the remainder. x^3+1/x-3

OpenStudy (anonymous):

i know the answer, i just need to know how to do it

jhonyy9 (jhonyy9):

x3 +1 : x-3 = x2 -3x -9 x3 -3x2 ------- 0 -3x2 +1 -3x2 +9x ---------- 0 -9x +1 -9x +27 --------- 0 -26

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

the reminaders actually 28

jhonyy9 (jhonyy9):

x3 +1 : x-3 = x2 -3x -9 and remainder -26 x3 +1 =(x-3)(x2 -3x -9) -26

OpenStudy (anonymous):

and the answer is x^2+3x+9

jhonyy9 (jhonyy9):

yes because i have missed one sign minus so but tha way how you can solve it is how i have wrote there

jhonyy9 (jhonyy9):

ok ?